1. Statistical measures of standalone risk Aa Aa Remember, the expected value of a probability distribution is a statistical measure of the average (mean) value expected to occur during all possible circumstances. To compute an assets expected return under a range of possible circumstances (or states of nature), multiply the anticipated return expected to result during each state of nature by its probability of occurrence. Consider the following case: Juan owns a two-stock portfolio that invests in Falcon Freight Company (FF) and Pheasant Pharmaceuticals (PP) Three-quarters of Juans portfolio value consists of FFs shares, and the balance consists of PPs shares. Each stocks expected return for the next year will depend on forecasted market conditions. The expected returns from the stocks in different market conditions are detailed in the following table Market Condition Probability of Occurrence Falcon Freight Pheasant Pharmaceuticals Strong Normal Weak 2090 35% 4590 3890 23% 30% 53% 30% 3890 Calculate expected returns for the individual stocks in Juans portfolio as well as the expected rate of return of the entire portfolio over the three possible market conditions next year. The expected rate of return on Falcon Freights stock over the next year is The expected rate of return on Pheasant Pharmaceuticalss stock over the next year is The expected rate of return on Juans portfolio over the next year isThe expected returns for Juans portfolio were calculated based on three possible conditions in the market. Such conditions will vary from time to time, and for each condition there will be a specific outcome. These probabilities and outcomes can be represented in the form of a continuous probability distribution graph. For example, the continuous probability distributions of rates of return on stocks for two different companies are shown on the following graph: PROBABILITY DENSITY Company G Company H 40-200 20 40 60 RATE OF RETURN (Percent] Based on the graphs information, which companys returns exhibit the greater risk? O Company H O Company G
Answer #1

1. Expected rate of return on FF's stock over the next year = Expected rate of return in strong market condition + Expected rate of return in normal market condition + Expected rate of return in weak market condition

Expected rate of return on FF's stock in strong market condition = P(Occurrence of Strong market condition) * (FF's return given strong market condition) = 20% * 38% = 0.2 * 0.38 = 0.076 or 7.6%

Expected rate of return on FF's stock in normal market condition = P(Occurrence of Normal market condition) * (FF's return given normal market condition) = 35% * 23% = 0.35 * 0.23 = 0.0805 or 8.05%

Expected market rate of return on FF's stock in weak market condition = P(Occurrence of Weak market condition) * (FF's return given weak market condition) = 45% * -30% = 0.45 * -0.30 = -0.135 or -13.5%

Therefore, expected rate of return on FF's stock over the next year = 7.6% + 8.05% - 13.5% = 2.15%

2. Expected rate of return on PP's stock over the next year = Expected rate of return in strong market condition + Expected rate of return in normal market condition + Expected rate of return in weak market condition

Expected rate of return on PP's stock in strong market condition = P(Occurrence of Strong market condition) * (PP's return given strong market condition) = 20% * 53% = 0.2 * 0.53 = 0.106 or 10.6%

Expected rate of return on PP's stock in normal market condition = P(Occurrence of Normal market condition) * (PP's return given normal market condition) = 35% * 30% = 0.35 * 0.30 = 0.105 0r 10.5%

Expected market rate of return on PP's stock in weak market condition = P(Occurrence of Weak market condition) * (PP's return given weak market condition) = 45% * -38% = 0.45 * -0.38 = -0.171 or -17.1%

Therefore, expected rate of return on PP's stock over the next year = 10.6% + 10.5% - 17.1% = 4%

3. Expected return of Juan's portfolio over the next year = (Weight of FF's stock in the portfolio * Expected rate of return on FF's stock over the next year) + (Weight of PP's stock in the portfolio * Expected rate of return on PP's stock over the next year)

= (0.75 * 2.15%) + (0.25 * 4%)

= 1.61% + 1%

= 2.61%

4. Risk comparison

Based on the continuous probability density graph of companies G and H, the probability as well as magnitude of negative rate of return is very low for company G as compared to that for company H. Hence, company H exhibits a greater risk.